TEST 2. PROBLEM SOLVING


Solve the problem below. Show your answers legibly, concisely, and completely. Use the back portion if necessary.


1. Determine the limiting reactant and excess reactant, when 36g of Al and 28.5g of HBr are reacted and how many grams of H_{2} gas are formed?

Al + HBr -> AlB*r_{3} + H_{2}



Given:


Find:


Solution:


1
Expert's answer
2022-04-11T10:26:05-0400

Given:

Mass of Al = 36 g

Mass of HBr = 28.5 g


Find:

The limiting reactant is ???

The excess reactant is ???

Mass of H2 = ???


Solution:

The molar mass of Al is 26.98 g/mol

The molar mass of HBr is 80.91 g/mol


Calculate the moles of each reactant:

Moles of Al = (36 g Al) × (1 mol Al / 26.98 g Al) = 1.334 mol Al

Moles of HBr = (28.5 g HBr) × (1 mol HBr / 80.91 g HBr) = 0.352 mol HBr


Balanced chemical equation:

2Al + 6HBr → 2AlBr3 + 3H2

According to stoichiometry:

2 mol of Al react with 6 mol of HBr

Thus, 1.334 mol of Al react with:

(1.334 mol Al) × (6 mol HBr / 2 mol Al) = 4.002 mol HBr

However, initially there is 0.352 mol of HBr (according to the task)

Therefore, HBr acts as limiting reactant and Al is excess reactant


According to stoichiometry:

6 mol of HBr produce 3 mol H2

Thus, 0.352 mol HBr produce:

(0.352 mol HBr) × (3 mol H2 / 6 mol HBr) = 0.176 mol H2


The molar mass of H2 is 2.016 g/mol

Therefore,

Mass of H2 = (0.176 mol H2) × (2.016 g H2 / 1 mol H2) = 0.355 g

Mass of H2 = 0.355 g


Answer:

The limiting reactant is HBr

The excess reactant is Al

0.355 grams of H2 gas are formed